OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 102 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to −6, how many women auditioned?

OpenStudy (tennistar):

The equation translated would need men to equal 102 - w, where w = # women. The equation translated would be [2(102-w)] - 3w = -6. (204 - 2w) - 3w 204 - 2w - 3w 204 - 5w = -6 -5w = -210 w = 42, there would be 60 men and 42 women.

OpenStudy (tennistar):

So your answer is 42

OpenStudy (anonymous):

Thank You!! Can you help me some more please?

OpenStudy (tennistar):

Sure

OpenStudy (anonymous):

ill post it on a new posting